Whole milk is 4% butterfat. How much skim milk with 0% butterfat should be added to 32 ounces of whole milk to obtain a mixture that is 2.5% butterfat?
My work thus far: oz btfat total
whole 32 4% 32(.04)

skim x 0% 0.00X
mix 32+x 2.5% 32+x(o.o25)
I don't know how to put the equation together but my answer is 19.2 oz

Answer by mananth(16949)

You are right
--------percent ---------------- quantity
whole milk 4 ---------------- 32 ounces
skim milk 0 ---------------- x ounces
Mixture 2.5 ---------------- 32 + x ounces

4*32+0x =2.5(32 +x)
128+0x= 80+2.5 x
0 x + -2.5 x = 80 - -128
-2.5 x = -48
/ -2.5
x = 19.2
19.2 ounces of 0 % fat skim milk